Numerical Analysis I, Fall 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 15

Assigned Dec 25, 2009. Due Jan 05, 2010.

1. Section 6.5: Problems 1, 2, 3, 5, 6, 7.

2. Section 6.6: Problems 5, 6, 7, 9, 10, 12, 14. For problem 14, read section 6.6.3.

Extra credit challenges: later.